How many liters of 20% and 60% acids must be mixed in order to make 12 liters of a 55% acid solution
Answer by stanbon(75887)

How many liters of 20% and 60% acids must be mixed in order to make 12 liters of a 55% acid solution
--------
Equation:
acid + acid = acid
0.20x + 0.60(12-x) = 0.55*12
-----
20x + 60*12-60x = 55*12
-----
-40x = -5*12
x = 12/8
---
x = 3/2 liters (amt. of 20% solution needed)
=================
Cheers,
Stan H.
=================